OpenStudy (anonymous):

What is the answwe to 27+4y you have to evaluate each expression

OpenStudy (anonymous):

what does y represent?

OpenStudy (jmark):

If 27+4y = 0, then we can evaluate the value of y Subtract 27 from both the sides 27+4y - 27 = -27 4y = -27 To isolate y, divide each side by 4 4y/4 = -27/4 y = -27/4 CBSE Class 7 Social Science Question Paper - http://education-help-online.blogspot.in/2014/08/cbse-class-7-social-science-question.html